# Quantum information with quantum-like bits

### v5

@@ IV.2 Connection to classical synchronization
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{X}}=\begin{pmatrix}A_{1}&-C\\ -C^{\top}&A_{2}\end{pmatrix}.$ (IV.13)
We show in panel a) of ??, the synchronization pattern in the dynamics of the Kuramoto model with correlation matrix $\mathcal{R}_{\mathrm{X}}$. In the transformation ??, the leading emergent eigenvector, associated to the largest eigenvalue $k+l$, is mapped from $\psi_{\uparrow}$ to $\psi_{\downarrow}$ [see ??]. The change of sign in the eigenvectors corresponds to a phase shift of $\Delta\phi=\pi$ between the first and the second half of the oscillators, as shown in the figure. In the case of the Hadamard gate, resources are transformed according to
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{H}}=\begin{pmatrix}A_{1}+A_{2}+C+C^{\top}&A_{1}-A_{2}-C+C^{\top}\\ A_{1}-A_{2}+C-C^{\top}&A_{1}+A_{2}-C-C^{\top}\end{pmatrix}.$ (IV.14)
The sum and difference of the basis graphs $A_{1}$ and $A_{2}$ in ?? highlight the fundamental role of this gate in creating a superposition state. The structure of ?? encodes key information about the expected synchronization patterns in the dynamics of the corresponding Kuramoto model, as depicted in panel b) of ??. Dephasing emerges between the first and second halves of the oscillators, primarily driven by cancellation effects in the two off-diagonal blocks, where the term $A_{1}-A_{2}$ plays a dominant role. When $A_{1}\simeq A_{2}$ and $C\simeq C^{\top}$, the off-diagonal blocks approach zero, leading to completely uncorrelated phases between the two oscillator subgroups over extended timescales. Furthermore, the second half of the oscillators, labeled by indices $i\in\{{N_{G}},\cdots 2{N_{G}}\}$ exhibits weaker synchronization compared to the first one. This is due to the fact that the matrix norm of the lower diagonal block in ?? is, on average, smaller than that of the upper diagonal block. Consequently, the second group of oscillators exhibits weaker correlations compared to the first group.
In this section, we derived general principles for mapping network resources in a manner consistent with an arbitrary set of QL gates. Additionally, we presented numerical evidence supporting the connection between QL gates and the steady-state synchronization of classical oscillator resources. In the following ??, we explore specific examples of experimental resources that could be utilized to implement our formalism in practice.
@@ V Experimental perspectives
So far, we have shown that the emergent states of synchronizing classical systems can be used as a basis to map the Hilbert space of a system of qubits. We discussed the primary role of emergent states in synchronization dynamics and the robustness of the latter in the presence of noise. In this section, we explore perspectives on implementing the proposed formalism using experimental physical resources. Additionally, we provide a high-level outline of computational and experimental validations that could support the further development of this framework.
For genuine quantum resources, the primary source of errors during information processing and measurement arises from uncontrolled noise. This noise can result from several factors, including external electromagnetic fields, unwanted interactions between qubits, fabrication defects, and imperfections in control protocols. Hou et al. (2016); Bengtsson et al. (2024); Funcke et al. (2022); Gupta and Nene (2021); Resch and Karpuzcu (2021) Addressing such errors typically requires multiple measurements to accurately resolve the final state of a quantum register after computations. This can introduce significant overhead in computing protocols and hinder the scalability of the approach to larger size. Error mitigation and decoding techniques, although conceptually scalable, can result in further limitations in size and performance. Bravyi et al. (2021); Ali et al. (2024) In our approach, error correction in the presence of noise translates to enhancing the robustness of long-time classical synchronization [see the discussion on panel b) of ??]. This ensures distinct and readily measurable transformations of the steady-state dynamics. Optimal synchronization can be achieved by numerically analyzing phase diagrams of synchronization indicators. For instance, in the case of the Kuramoto model, time-resolved order parameters offer clear insight into the synchronization strength. Basnarkov and Urumov (2007); Wang, Xu, and Zheng (2021) Parameter domains with the highest degree of synchronization can be identified as optimal candidates for experimental implementation.
@@
Implementing coupled LC circuits is a straightforward approach to implement macroscopic experimental resources that exhibit classical synchronization. Allam et al. (2006); Barbi (2021) Numerical analyses from Ref. Maffezzoni et al., 2015 show that this class of systems can be effectively approximated by Kuramoto-type models for medium-scale systems of 60 oscillators. The couplings between the model oscillators are tuned by adjusting the transconductance values of the circuit transistors.
@@
Spin-torque oscillators (STO) emerge as another promising candidate for implementing QL information processing. Raychowdhury et al. (2019) STOs are nonlinear magnetic oscillators that can be integrated into nanoscale devices operating at microwave frequencies. They have the ability to synchronize in both frequency and phase with external oscillatory signals as well as with other STOs. The full magnetization dynamics of these systems, defined by the Landau-Lifshitz-Gilbert-Slonzewski (LLGS) equation, can be simplified to the Kuramoto model for small-amplitude oscillations. Belanovsky et al. (2013); Kanao et al. (2019) Ref. Flovik, Macià, and Wahlström, 2016 demonstrates the experimental feasibility of modeling synchronizing classical systems with up to a few hundred oscillators. Beyond this scale, technical constraints arising from finite-size effects restrict the extent of achievable synchronization.
@@
Metal-to-insulator transition (MIT) devices leverage materials that can switch between insulating and metallic (conductive) states in response to external conditions such as temperature, pressure, or electrical stimuli. Kim et al. (2010) Research on small-scale MIT circuits has demonstrated their fundamental ability to produce classical synchronization behavior through voltage-triggered oscillatory dynamics. Kim et al. (2010); Parihar et al. (2015); Raychowdhury et al. (2019) One limitation of current MIT devices is their low oscillation frequency, typically below 1 MHz, which hinders the ability to achieve fine-resolution synchronization and detect subtle variations in coupled oscillator behavior. However, ongoing advancements in scaling down these devices are expected to increase their oscillation frequencies, enabling more precise and time-resolved synchronization measurements. Cotter et al. (2014)
@@
Building on the experimental perspectives outlined above, it appears feasible to validate our QL-bit theory using networks of up to a hundred oscillators. Such an experimental setup would enable the implementation of proof-of-concept algorithms on systems with ${N_{\mathrm{QL}}}=2$ or ${N_{\mathrm{QL}}}=3$ QL-bits, where each QL-bit comprises a small network of 555510101010 oscillators. A particularly intriguing class of quantum information processes for systems of this scale includes those that generate entanglement, such as the formation of Bell states (for ${N_{\mathrm{QL}}}=2$) and their higher-order generalizations, the Greenberger–Horne–Zeilinger (GHZ) states. Greenberger, Horne, and Zeilinger (1989); Erhard, Krenn, and Zeilinger (2020) This analysis has the potential to provide valuable insights into fundamental questions surrounding the classical-to-quantum correspondence. In particular, it could help clarify the role of non-emergent states in classical synchronizing systems as potential hidden variables within a QL-based framework for analyzing entanglement. Scholes (2025b)
We conclude this section by presenting a practical perspective on implementing an experimental scheme to test our approach:
@@
    Engineer a classical synchronizing system: Design a system with two-body correlations represented by a graph, ensuring that its leading emergent state corresponds to the desired initial state of a QL algorithm. For example, the factorized state $\psi_{\uparrow}$ can be associated with a Cartesian-product network structured as described in ??.
@@
    Initialize oscillator phases: Assign arbitrary initial conditions to the oscillator phases, ensuring they adhere to the constraint that the system’s long-time steady state achieves the expected synchronization.
@@
    Allow the classical dynamics to equilibrate: Let the system evolve until it reaches a steady state. This step requires time-resolved measurements to ensure the system has sufficient time to reach a long-time synchronized state.
@@
    Perform gate operations: Each gate operation consists of two phases: (a) transforming the correlations between the degrees of freedom and (b) allowing the system to synchronize and relax to a new emergent state.
  5. 5.
    Measure the final state: Apply projection operators to extract the state of the final resource after all gate operations have been implemented.
The proposed scheme can be tested numerically by measuring multi-time correlation functions. This process involves a series of time evolution operators defined by gate-transformed resources at different stages of a given QL algorithm. Preliminary analytical work on the Kuramoto dynamics is discussed in ?? [see ??].
@@ VI.1 Axiomatic definition
and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$ [??]. The QL measurement of $\mathcal{F}$ on a state $\psi$ is defined by the map $\psi\mapsto\mathcal{F}\psi$.
@@
the QL probability of collapsing onto the subspace spanned by the eigenvector $f_{\bm{\sigma}}$ with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$. ?? is expressed in terms of the projector
$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}\otimes f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (VI.4)
@@
In particular, let us consider two generic QL observable operators $\mathcal{F}$ and $\mathcal{G}$ whose spectra are defined by
$\mathcal{F}f_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}},\hskip 28.45274pt\mathcal{G}g_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}.$ (VI.5)
For a nonzero state $\psi$ we denote by
@@
We show in the next ?? that the QL probability model defined via Axioms 1 to 5 violates by construction Kolmogorovian classical statistics. This feature allows to introduce a notion of QL interference in our approach. Khrennikov (2003, 2019)
@@ VI.2 Quantum probabilities in a QL model
Before introducing a formula for the update of QL probabilities, it is instructive to briefly recall its classical correspondent, to assess analogies and differences between the two. A classical probability space is defined by the triplet $\{\Omega,\Sigma,P^{(\mathrm{cl})}\}$, where $\Omega$ is a space of experimental outcomes, $\Sigma$ the $\sigma$-algebra of its subsets, and $P^{(\mathrm{cl})}$ a probability measure on $\Sigma$.Shiryaev and Wilson (2013) Be $\{\mathcal{A}_{j}\}_{j=1}^{J}$ a partition of $\Omega$, i.e. a collection of $J$ sets defined such that $A_{j}\neq\emptyset\;\forall j$, $\mathcal{A}_{j}\cap A_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ and $\cup_{j=1}^{J}\mathcal{A}_{j}=\Omega$. $P^{(\mathrm{cl})}$ is a positive measure on the partition elements, that is $P^{(\mathrm{cl})}(\mathcal{A}_{j})>0\;\forall j$, and it is normalized: $\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{A}_{j})=1$. For every $\mathcal{B}\in\Sigma$ the classical formula of total probabilities holds: Kolmogoroff (1933)
@@
We discuss now the correspondent of ?? for the present QL framework. We restrict the following analysis to couples of non-degenerate QL observables, $\mathcal{F}$ and $\mathcal{G}$, whose eigenvalue problems are defined as in ??. The eigenvalues of these operators are associated with a one-dimensional subspaces, which can be fixed by selecting any normalized eigenvector. In this case, the QL conditional probabilities defined in ?? lose their dependence on the initial state $\psi$, and can be simply expressed as
@@
An extended derivation of ?? is provided in ??, where we give the explicit expression of the nonclassical interference coefficient $\Lambda^{\psi}_{\bm{\sigma},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}}$ [??]. Information processing can be designed in a QL way, to enhance or suppress some spectral measurements compared to classical computing protocols. Khrennikov (2021)
As noted in Ref. Khrennikov, 2021, a significant feature of the QL probability formula ?? is that it bypasses the need for calculating joint probability distributions, which are typically exponential in computational complexity. In contrast, such calculations are required for the classical updates outlined in ??. This distinction relates directly to the foundational principles of quantum advantage. We intend to further investigate this property in future testing of QL algorithms.

### v6

@@ IV.2 Connection to classical synchronization
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{X}}=\begin{pmatrix}A_{1}&-C\\ -C^{\top}&A_{2}\end{pmatrix}.$ (IV.14)
We show in panel a) of ?? the synchronization pattern in the dynamics of the Kuramoto model with correlation matrix $\mathcal{R}_{\mathrm{X}}$. In the transformation ??, the leading emergent eigenvector, associated to the largest eigenvalue $k+l$, is mapped from $\psi_{\uparrow}$ to $\psi_{\downarrow}$ [see ??]. The change of sign in the eigenvectors corresponds to a phase shift of $\Delta\phi=\pi$ between the first and the second half of the oscillators, as shown in the figure. In the case of the Hadamard gate, resources are transformed according to
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{H}}=\begin{pmatrix}A_{1}+A_{2}+C+C^{\top}&A_{1}-A_{2}-C+C^{\top}\\ A_{1}-A_{2}+C-C^{\top}&A_{1}+A_{2}-C-C^{\top}\end{pmatrix}.$ (IV.15)
The structure of ?? encodes information about the expected synchronization patterns in the dynamics of the corresponding Kuramoto model, as depicted in panel b) of ??. Dephasing emerges between the first and second halves of the oscillators, primarily driven by cancellation effects in the two off-diagonal blocks. When $A_{1}\approx A_{2}$ and $C\approx C^{\top}$, these blocks approach zero, decorrelating the phases between the two oscillator subgroups. Furthermore, the second half of the oscillators, labeled by indices $i\in\{{N_{G}},\ldots,2{N_{G}}\}$ exhibits weaker synchronization compared to the first half. This is due to the fact that the matrix norm of the lower diagonal block in ?? is, on average, smaller than that of the upper diagonal block. Consequently, the second group of oscillators exhibits weaker correlations compared to the first group. The interested reader is referred to Ref. Amati and Scholes, 2025. There, we show with a straightforward generalization of the present Kuramoto model that this kind classical synchronization patterns can accurately represent arbitrary quantum states.
In this section, we derived general principles for mapping network resources in a manner consistent with an arbitrary set of QL gates. Additionally, we presented numerical evidence supporting the connection between QL gates and the steady-state synchronization of classical oscillator resources. In the following ??, we explore specific examples of experimental resources that could be utilized to implement this approach in practice.
@@ V Experimental perspectives
So far, we have shown that the emergent states of synchronizing classical systems can be used as a basis to map the Hilbert space of a system of qubits. We discussed the primary role of emergent states in synchronization dynamics and the robustness of the latter in the presence of noise. Additionally, we designed a general approach to implement gate transformations on the network resources. In this section, we explore perspectives on implementing the proposed formalism using experimental physical resources. Additionally, we provide a high-level outline of computational and experimental validations that could support the further development of this framework.
For genuine quantum resources, the primary source of errors during information processing and measurement arises from uncontrolled noise. This noise can result from several factors, including external electromagnetic fields, unwanted interactions between qubits, fabrication defects, and imperfections in control protocols. Hou et al. (2016); Bengtsson et al. (2024); Funcke et al. (2022); Gupta and Nene (2021); Resch and Karpuzcu (2021) Addressing such errors typically requires multiple measurements to accurately resolve the final state of a quantum register after computations. This can introduce significant overhead in computing protocols and hinder the scalability of the approach to larger size. Error mitigation and decoding techniques, although conceptually scalable, can result in further limitations in size and performance. Bravyi et al. (2021); Ali et al. (2024) In our approach, error correction in the presence of noise could be achieved by encoding information into a larger number of oscillators. Given that those are defined on a classical phase space, the related cost of the strategy scales polynomially with the number of resources. This suggests a simple strategy to implement a QL version of repetition codes in our approach.
@@
Implementing coupled LC circuits is a straightforward approach to build macroscopic resources that exhibit classical synchronization. Allam et al. (2006); Barbi (2021) Numerical analyses from Ref. Maffezzoni et al., 2015 show that this class of systems can be effectively approximated by Kuramoto-type models for medium-scale systems of 60 oscillators. The couplings between the model oscillators are tuned by adjusting the transconductance of the differential pair transistors.
@@
Spin-torque oscillators (STO) emerge as another promising candidate for implementing QL information processing. Raychowdhury et al. (2019) STOs are nonlinear magnetic oscillators that can be integrated into nanoscale devices operating at microwave frequencies. They have the ability to synchronize in both frequency and phase with external oscillatory signals as well as with other STOs. The full magnetization dynamics of these systems, defined by the Landau-Lifshitz-Gilbert-Slonzewski equation, can be simplified to the Kuramoto model for small-amplitude oscillations. Belanovsky et al. (2013); Kanao et al. (2019) Ref. Flovik, Macià, and Wahlström, 2016 demonstrates the experimental feasibility of modeling synchronizing classical systems with up to a few hundred STOs. Beyond this scale, technical constraints arising from finite-size effects restrict the extent of achievable synchronization.
@@
Metal-to-insulator transition (MIT) devices leverage materials that can switch between insulating and metallic (conductive) states in response to external conditions such as temperature, pressure, or electrical stimuli. Kim et al. (2010) Research on small-scale MIT circuits has demonstrated their fundamental ability to produce classical synchronization behavior through voltage-triggered oscillatory dynamics. Kim et al. (2010); Parihar et al. (2015); Raychowdhury et al. (2019) One limitation of current MIT devices is their low oscillation frequency, typically below 1 MHz, which hinders the ability to achieve fine-resolution synchronization and detect subtle variations in coupled oscillator behavior. However, ongoing advancements in scaling down these devices are expected to increase their oscillation frequencies, enabling more accurate time-resolved synchronization measurements. Cotter et al. (2014)
@@
Building on the perspectives discussed above, experimental validation of our theory using networks of up to a hundred oscillators appears feasible. In such a setup, each QL bit could consist of a small network of 5–10 oscillators, allowing for proof-of-concept implementations with ${N_{\mathrm{QL}}}=2$ or ${N_{\mathrm{QL}}}=3$ QL bits. At this scale, a particularly interesting class of quantum information processes includes entanglement generation, such as the formation of Bell states (for ${N_{\mathrm{QL}}}=2$) and their higher-order generalizations, the Greenberger–Horne–Zeilinger (GHZ) states. Greenberger, Horne, and Zeilinger (1989); Erhard, Krenn, and Zeilinger (2020) Advancements in this framework are given in Ref. Amati and Scholes, 2025.
Exploring small-scale quantum processes might offer a possible route to address some foundational questions. For example: To what extent can one construct a classical analogue of entanglement in a classical many-body system? Is there a fundamental limit beyond which even infinite classical resources fail to replicate essential features of entanglement? Such a limitation could signal an intrinsic boundary between which kind of quantum behaviors can be reproduced by classical dynamics. We argue that our approach can serve as a platform to investigate these general problems.
We end this section with a practical outlook on how our approach could be tested experimentally.
@@
    Engineer a classical synchronizing system: Design a classical system of nonlinear oscillators with two-body interactions, defined by a graph. This graph should be constructed so that its leading emergent state matches the desired initial state of a quantum circuit. For example, the factorized state $\psi_{\uparrow}$ corresponds to the eigenvector with the largest eigenvalue of a Cartesian-product network, as given in ??. The other parameters of the classical system should be tuned to ensure the emergence of synchronization at long times.
@@
    Transform the network according to arbitrary quantum logic: Apply to the network a unitary transformation corresponding to a given quantum circuit, following the protocol proposed in ??.
@@
    Allow the classical dynamics to equilibrate: Assign random initial phases to the classical oscillators and let the system evolve until it reaches a synchronized steady state. We discuss formally in Ref. Amati and Scholes, 2025 how this steady states encodes information on the gate initially applied to the classical network.
@@
    Measure the final state: Calculate the overlap between the oscillators’ steady state and the eigenstates of arbitrary observable operators. The resulting projection coefficients can then be interpreted as nonclassical probability amplitudes within the framework of QL theories, as discussed in detail in the next ??.
In the next and final ?? we expand over this last point, to show that our approach allows to mimic non-Kolmogorov statistics, that violates the Bayes theorem due to the occurrence of nonclassical interference.
@@ VI.1 Axiomatic definition
and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$ [??].
@@
the QL probability of collapsing onto the subspace spanned by the normalized eigenvector $f_{\bm{\sigma}}$ with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$. ?? is expressed in terms of the projector
$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (VI.4)
@@
In particular, let us consider two generic QL observable operators $\mathcal{F}$ and $\mathcal{G}$ whose spectra decompositions are defined by
$\mathcal{F}=\sum_{\bm{\sigma}\in{\bm{\mathcal{S}}}}\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}}f_{\bm{\sigma}}^{\dagger},\hskip 28.45274pt\mathcal{G}=\sum_{\bm{\sigma}\in{\bm{\mathcal{S}}}}\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}g_{\bm{\sigma}}^{\dagger},$ (VI.5)
respectively. For a nonzero state $\psi$ we denote by
@@
We show in the next ?? that the stochastic model defined via Axioms 1 to 5 violates by construction Kolmogorovian classical statistics. This feature allows to introduce a notion of QL interference in our approach. Khrennikov (2003, 2019)
@@ VI.2 Quantum probabilities in a QL model
Before introducing a formula for the update of QL probabilities, it is useful to briefly recall its classical correspondent, to assess analogies and differences between the two. A classical probability space is defined by the triplet $\{\Omega,\Sigma,P^{(\mathrm{cl})}\}$, where $\Omega$ is a space of experimental outcomes, $\Sigma$ the $\sigma$-algebra of its subsets, and $P^{(\mathrm{cl})}$ a probability measure on $\Sigma$.Shiryaev and Wilson (2013) Be $\{\mathcal{A}_{j}\}_{j=1}^{J}$ a partition of $\Omega$, i.e. a collection of $J$ sets defined such that $\mathcal{A}_{j}\neq\emptyset\;\forall j$, $\mathcal{A}_{j}\cap A_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ and $\cup_{j=1}^{J}\mathcal{A}_{j}=\Omega$. $P^{(\mathrm{cl})}$ is a positive measure on the partition elements, that is $P^{(\mathrm{cl})}(\mathcal{A}_{j})>0\;\forall j$, and it is normalized: $\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{A}_{j})=1$. For every $\mathcal{B}\in\Sigma$ the classical formula of total probabilities holds: Kolmogoroff (1933)
@@
We discuss now the correspondent of ?? for the present QL framework. We restrict the following analysis to couples of non-degenerate observables, $\mathcal{F}$ and $\mathcal{G}$, whose eigenvalue problems are defined as in ??. The eigenvalues of these operators are associated with one-dimensional subspaces, which can be fixed by selecting any eigenvector. In this case, the QL conditional probabilities defined in ?? lose their dependence on the initial state $\psi$, and can be simply expressed as
@@
An extended derivation of ?? is provided in ??, where we give the explicit expression of the nonclassical interference coefficient $\Lambda^{\psi}_{\bm{\sigma},\bm{\sigma}^{\prime},\bm{\sigma}^{\prime\prime}}$ [??].
We conclude this section by connecting the abstract discussion above to an operational strategy to simulate interference phenomena. In Ref. Amati and Scholes, 2025 we prove that the total phase vector of a system of classical oscillators can be expanded over the sum of all eigenvalues of the graph, including both emergent contributions (isolated peaks) and non-emergent terms. However, only the emergent states are associated to a quantum Hilbert space. To resolve the quantum information encoded in the system it is therefore required to maximize the spectral gap between emergent and random states. This is achieved via an ad-hoc choice of the valencies of the elementary graphs comprising the network, as discussed in ??. Conveniently, this preparation stage needs to be defined once for all for any circuit, as the graph spectra are preserved by unitary QL gates. Let consider as an example the specific case of LC circuits, already introduced in ??. It is shown in Ref. Maffezzoni et al., 2015 that the oscillatory patterns of these circuits allow to build Kuramoto-type models, where the interactions between oscillators are defined from transconductance elements. After building this network and allowing for its long-time synchronization, a readout of the phases of the oscillators phases is required. As the voltage $V_{j}(t)$ of each LC oscillator is related to its angle via $V_{j}(t)\propto\cos\left[\theta_{j}(t)\right]$, classical phase readout in this setup can be implemented by means of voltage probes, such as oscilloscopes or lock-in amplifiers. Finally, QL measurement processes can be simulated by numerically postprocessing the classical phase vector. For example, conditional probablities from ?? can be calculated by projecting the phases of the oscillators onto the spectrum of a QL observable operator $\mathcal{F}$. These nonclassical probabilities obey by construction the five axioms listed above, mimicking nonclassical interference phenomena.